OpenStudy (anonymous):

Two semi trucks are driving loads from Chicago to Denver, distance of 1125 miles. First truck leaves at 8 AM, averages 50 mph. Second truck leaves at 9 AM, averages at 54 mph. a. How far are they from Chicago when the second truck overtakes the first truck? b. If each of the drivers stop for meals for a total of 3 hours, what time will it be in Chicago when each truck reaches Denver? c. How much would the driver of the first truck have to increase her speed in order to arrive in Denver first?

OpenStudy (anonymous):

I already found the two parameter equations: 50 t and 54(t-1) ... but now I don't know what to do. The answers to the questions are a) 675 miles; b) 9:30 AM, 8:50 AM; c) at least 1.53 mph . But I don't know how to get to these.

OpenStudy (anonymous):

do you have to explain how u got your answer or is it multiple choice ?

OpenStudy (anonymous):

I have to show my work

OpenStudy (anonymous):

your equations are wrong slightly first truck goes 50 miles before second so equation is 50(t+1) second is 54t equate t=50/4 distance =vel *time =50/4 *54 =675

OpenStudy (anonymous):

how'd you get t=50/4?

OpenStudy (anonymous):

(where'd the 4 come from?)

OpenStudy (anonymous):

50(t+1)=54t 50t+50=54t 50=54t-50t=4t t=50/4

OpenStudy (anonymous):

Okay thanks. How about part b?

OpenStudy (anonymous):

take truck 2 distance/speed =1125/54 =20.83 = 20 hrs 50 min +3 hrs stop time =23 hrs 50 min add on when he left 9am =almost a whole day ie 8. 50 you do the rest

OpenStudy (anonymous):

thats just for the second truck right?

OpenStudy (anonymous):

okay I got the first truck's time

OpenStudy (anonymous):

how about the last part? (part c)

OpenStudy (anonymous):

@jamaica25

OpenStudy (anonymous):

if truck (1) got the 50 mile start he would have 1075 miles still to go truck (2) takes 20.83 hrs so truck(1) would would have to do 1075/ 20.83 hrs=52